# Risk-Aware Autonomous Driving with Linear Temporal Logic Specifications

### v2

@@ IV-B Implementation Framework
The above discussion is based on a MDP model of vehicles. However, to deploy this approach in a real car, we need to abstract an MDP model from vehicle dynamics, as well as determine how to implement high-level policies obtained by solving the LP problem. The complete framework for implementation on autonomous vehicles is illustrated in Fig. 5. The related codes are available in [36].
@@
Vehicle dynamics: In the experiments, we consider the dynamics of vehicles with a discrete-time bicycle model. Specifically, the system state consists of the horizontal position, heading angle, and linear speed of the vehicle, denoted by $[p_{x},p_{y}]$, $\theta$, and $v$, respectively. The control inputs to the vehicle are steering angle and acceleration, denoted by $\phi$ and $a$. Consequently, the system state is $\mathbf{x}(t)\!=\![p_{x}(t),p_{y}(t),\theta(t),v(t)]^{T}\!\in\!\mathbb{X}\!\subset\!\mathbb{R}^{4}$, and the system input is $\mathbf{u}(t)=[\phi(t),a(t)]^{T}\!\in\!\mathbb{U}\!\subset\!\mathbb{R}^{2}$, where $\mathbb{X}$ and $\mathbb{U}$ are state space and input space, respectively. The kinematic model of the vehicle is a bicycle model $\mathbf{x}(t+1)\!=\!f(\mathbf{x}(t),\mathbf{u}(t))$ of which the details can be referred to in [37]. To capture the environmental uncertainty, a zero-mean Gaussian-distributed disturbance vector, $\mathbf{w}\sim\mathcal{N}(0,\Sigma)$, is considered with the variance $\Sigma\in\mathbb{R}^{4\times 4}$. The vehicle model with uncertainty is denoted by $\mathbf{x}(t+1)\!=\!f(\mathbf{x}(t),\mathbf{u}(t))+\mathbf{w}(t)$.
From continuous dynamics to MDP: To abstract $\mathcal{M}_{v}$ from the vehicle dynamics model with uncertainty, we first discretize the state space $\mathbb{X}$ by gridding, where the finite state set $S$ comprises the centers of the grid cells. For simplicity, this implementation considers only the horizontal positions of vehicles in $\mathcal{M}_{v}$. Similarly, the continuous input space $\mathbb{U}$ is discretized into a finite action set $A$ using appropriate interface functions. In this implementation, the transition probabilities are predefined by experts, which can also be generated autonomously and formally [19]. Additionally, the set of atomic propositions $\mathsf{AP}$ is determined by corresponding propositions in specifications, and the labeling function $\mathcal{L}$ can be extracted from the semantic maps in the perception module. Since perception is beyond the scope of this paper, the labels are manually assigned in this implementation.
@@
where $N$ denotes a horizon number, $X_{0}$ is the current state, $\mathbf{Q}\!\in\!\mathbb{R}^{4\times 4}$ and $\mathbf{R}\!\in\!\mathbb{R}^{2\times 2}$ are diagonal weight matrixes. The first optimal control input is deployed to the vehicle.
@@ IV-C Pedestrian Crossing Example
We first consider a simple scenario with a dynamic environment where a pedestrian is crossing at a crosswalk. The ego vehicle is expected to respond appropriately to the pedestrian’s crossing behavior. To this end, we build an MC model to represent the uncertain behavior of the pedestrian. The specification of the ego vehicle is $\square(p\!\rightarrow\!\neg\!c)\wedge\lozenge(t)$, where $c$ and $t$ denote the crosswalk and target zone respectively, and $p$ represents the presence of a pedestrian on the crosswalk. The cost of hitting a pedestrian is set to $8$. The ego vehicle is expected to autonomously decide whether to stop or proceed based on the real-time situation. From the risk curve shown in Fig. 6(a), the LP problem in Eq. (13) ensures that the risk always remains below the threshold. To test risk sensitivity, we set different risk thresholds $r_{th}$ for the LP problem. The simulation results, depicted in Fig. 6(b), indicate that the ego vehicle will stop at a smaller clearance ahead of the crosswalk as the increase of risk threshold, i.e., $r_{th}=0.1,1,5$. But $r_{th}=10$ is too high to make the ego car account for the collision risk such that the ego car directly hits the pedestrian, which is undesirable. This result demonstrates the parameter $r_{th}$ can be used to tune the conservatism of the synthesized policy.
(a) Risk curve for $r_{th}\!=\!1$
(b) Stopping points for different risk thresholds.
Figure 6: Pedestrian crossing scenario
In more complex scenarios, the LP problem encounters infeasibility in risky states, requiring a larger $r_{th}$. But a larger $r_{th}$ will lead to unreasonably aggressive behavior in less risky states. As a result, while the formulation in (13) performs well in the pedestrian scenario, it generates unreasonable behavior in complex scenarios, such as unprotected turns in Fig. 1, which cannot be resolved by simply adjusting $r_{th}$. To address this issue, we propose an improved method in the next section.
@@ V Improved risk-aware control and experimental study
Ideally, human drivers can actively minimize real-time risks while avoiding not only catastrophic outcomes but also overly conservative behavior. To achieve such a multi-objective balance, we introduce a soft risk threshold $r_{s}$ and a hard risk threshold $r_{h}$ to manage risk at different levels. This gives the following improved risk-aware control synthesis problem written as an LP optimization,
@@
where $\xi$ denotes the relaxation variable for soft risk threshold $r_{s}$ and $K\!>\!0$ is a constant to balance satisfying the co-safety specifications and minimizing the risks of safety specifications. The risk metric is minimized only when the target term forces the risk to exceed the soft threshold. Therefore, $r_{s}$ filters out overly conservative behavior, while $r_{h}$ ensures critical safety.
@@ V-A Experimental Study
We validate the improved risk-aware framework in three different scenarios, using default parameters: risk thresholds $r_{s}\!=\!1$ and $r_{h}\!=\!2$, and discounting factor $\gamma\!=\!0.8$. The related video demonstration is accessible at the link  [^1] https://youtu.be/r5kEMW8oPQE.
1) Pedestrian crossing scenario: The improved method ensures safety and generates behavior with tunable conservatism under the same settings as in Sec. IV-C, demonstrating that the improved LP problem encompasses functionalities of the original one in (13). The maximum and average risk values in the process are listed in Table I.
2) Unexpected construction scenario: Real-world traffic frequently encounters unexpected situations that make the original specifications infeasible [38], such as traffic participants violating rules or disruptions of traffic environments. We tested the proposed methods in such a scenario where an unexpected construction zone appears in the forward lane of the ego car, as illustrated in Fig. 7. In this case, the ego vehicle must deviate from a part of the original safety specifications to reach its target. The specification of this scenario for the ego vehicle is ($\square(\neg c\wedge\neg o\wedge\neg s)\wedge\lozenge(t)$, where $c$, $o$, $s$ and $t$ are labels for construction areas, opposite lanes, sidewalks, and target areas. We assign cost values of $5$, $3$, and $1$ to the construction area, sidewalk, and opposite lane, respectively. Figure 7 shows the generated trajectory, where the ego vehicle autonomously chooses to bypass the construction zone via the opposite lane with minimal violation and finally reaches the target. This result demonstrates that the improved control method effectively balances different risks and minimizes violations. This advantage arises from incorporating severity into LTL specifications.
@@
3) Unprotected turn scenario: We consider a complex intersection scenario that includes a traffic light system and dynamic opponent vehicles, as depicted in Fig. 1. The ego vehicle (blue car) is expected to follow traffic signals and avoid both non-drivable areas and the opponent vehicles (red car). The specification has been introduced in the study case of Sec. II-B. An MC model is designed to represent the switching of traffic lights and the movement of the opponent vehicle. For simplicity, we assume the opponent vehicle always moving forward. The collision risk with the opponent vehicle is evaluated by calculating the overlap between risk fields of vehicles, following the method described in [24]. With the policy synthesized by the LP problem, the ego vehicle will wait at the boundary of the intersection until the light turns green, adjust its decisions based on the real-time positions of the opponent vehicles, and finally reach the target position. The whole process is illustrated in Fig. 1 and the attached video. This case accounts for traffic lights, non-drivable areas, and other dynamic vehicles, which are typical factors in traffic. Thus, this method can handle complex situations and has the potential to scale to various scenarios.
To examine the influence of different discounting factors, the corresponding risk curves are shown in Fig. 8. While the generated behavior remains identical across different discounting factors, the risk curves in the process are different. Compared to $\gamma\!=\!0.8$, the discounting factor $\gamma\!=\!0.5$ places more emphasis on recent events, causing the risk curves of $\gamma\!=\!0.5$ to gradually increase as the two vehicles approach each other. In contrast, the risk curves for $\gamma\!=\!0.8$ remain flatter because the encounter is perceived earlier. Since placing more focus on recent events lowers perceived risk values, the thresholds are adjusted accordingly.
@@
Table I presents the maximum and average risk values with the default discounting and threshold, along with the state number and computation time for the LP problem in (13). The results show that the improved LP problem effectively manages different types of risk and maintains gross risk value within the predefined level throughout the execution process. However, computation time increases significantly as the number of states grows such that a coarse abstraction is adopted. Future work will explore adaptive abstraction and approximation methods to improve computational efficiency.
@@
TABLE I: Risk value and computation time in the three scenarios.
@@ VI Conclusion
In this paper, we extend a human-like risk metric to account for complex traffic scenarios by integrating LTL specifications. On the other hand, the proposed metric enhances the practicality of temporal logic in traffic tasks by incorporating human-like awareness. Furthermore, a risk-aware control method is improved to balance different types of risks in a human-like manner. Extensive simulation results validate the effectiveness and potential of this framework and illustrate the impact of various parameters. To better handle realistic traffic, future work should incorporate strategic interactions between vehicles, reduce computational costs for finer abstractions, and learn parameters from real-road data instead of manual configuration.

### v3

@@ IV-B Implementation Framework
The above discussion is based on an MDP model of vehicles. However, to deploy this approach in a real car, we need to abstract an MDP model from vehicle dynamics and determine how to implement high-level policies obtained by the LP problem. The overall framework is shown in Fig. 5.
@@
Vehicle dynamics: In the experiments, we consider the dynamics of vehicles with a discrete-time bicycle model. Specifically, the system state consists of the planar position, heading angle, and linear speed of the vehicle, denoted by $[p_{x},p_{y}]$, $\theta$, and $v$, respectively. The control inputs to the vehicle are steering angle and acceleration, denoted by $\phi$ and $a$. Consequently, the system state is $\mathbf{x}(t)\!=\![p_{x}(t),p_{y}(t),\theta(t),v(t)]^{T}\!\in\!\mathbb{X}\!\subset\!\mathbb{R}^{4}$, and the system input is $\mathbf{u}(t)=[\phi(t),a(t)]^{T}\!\in\!\mathbb{U}\!\subset\!\mathbb{R}^{2}$, where $\mathbb{X}$ and $\mathbb{U}$ are state space and input space, respectively. The kinematic model of the vehicle is a bicycle model $\mathbf{x}(t+1)\!=\!f(\mathbf{x}(t),\mathbf{u}(t))$, and its details can be found in [26]. To capture the environmental uncertainty, a zero-mean Gaussian-distributed disturbance vector, $\mathbf{w}\sim\mathcal{N}(0,\Sigma)$, is considered with the variance $\Sigma\in\mathbb{R}^{4\times 4}$. The vehicle model with uncertainty is denoted by $\mathbf{x}(t+1)\!=\!f(\mathbf{x}(t),\mathbf{u}(t))+\mathbf{w}(t)$.
From continuous dynamics to MDP: To abstract $\mathcal{M}_{v}$ from the vehicle dynamics model with uncertainty, we first discretize the state space $\mathbb{X}$ by gridding, where the finite state set $S$ comprises the centers of the grid cells. For simplicity, this implementation considers only the planar positions of vehicles in $\mathcal{M}_{v}$. Similarly, the continuous input space $\mathbb{U}$ is discretized into a finite action set $A$ using appropriate interface functions. In this implementation, the transition probabilities are predefined by experts, which can also be generated autonomously and formally [10]. Additionally, the set of atomic propositions $\mathsf{AP}$ is determined by corresponding propositions in specifications, and the labeling function $\mathcal{L}$ can be extracted from the semantic maps in the perception module. Since perception is beyond the scope of this paper, the labels are manually assigned in this implementation.
@@
where $N$ denotes a horizon number, $X_{0}$ is the current state, $\mathbf{Q}\!\in\!\mathbb{R}^{4\times 4}$ and $\mathbf{R}\!\in\!\mathbb{R}^{2\times 2}$ are diagonal weight matrices with positive weights.
@@ IV-C Pedestrian Crossing Example
We first consider a simple scenario with a dynamic environment where a pedestrian is crossing at a crosswalk. The ego vehicle is expected to respond appropriately to the pedestrian’s crossing behavior. To this end, we build an MC model to represent the uncertain behavior of the pedestrian. The specification of the ego vehicle is $\square(p\!\rightarrow\!\neg\!c)\wedge\lozenge(t)$, where $c$ and $t$ denote the crosswalk and target zone, respectively, and $p$ represents the presence of a pedestrian on the crosswalk. The cost of hitting a pedestrian is set to 888. The ego vehicle is expected to autonomously decide whether to stop or proceed based on the real-time situation. To test risk sensitivity, we set different risk thresholds $r_{th}$ for the LP problem. The simulation results, depicted in Fig. 6, indicate that the ego vehicle will stop at a smaller clearance ahead of the crosswalk as the risk threshold increases, i.e., $r_{th}=0.1,1,5$. However, when $r_{th}=10$, the collision risk is ignored, resulting in the ego vehicle striking the pedestrian. This result demonstrates that the parameter $r_{th}$ can be used to tune the conservatism of the synthesized policy.
Figure 6: Stopping points for different risk thresholds.
However, tuning $r_{th}$ becomes an issue in more complex scenarios. The LP problem might be infeasible in risky states, requiring a larger $r_{th}$, but increasing $r_{th}$ leads to unreasonably aggressive behavior in less risky states. As a result, while the formulation in (13) performs well in the pedestrian scenario, reasonable behavior cannot be achieved by tuning $r_{th}$ in complex scenarios such as unprotected turns in Fig. 1. To address this issue, we propose an improved method in the next section.
@@ V Improved risk-aware control and experimental study
Ideally, human drivers can actively minimize real-time risks while avoiding not only catastrophic outcomes but also overly conservative behavior. To achieve such a multi-objective balance, we introduce a soft risk threshold $r_{s}$ and a hard risk threshold $r_{h}$ to manage risk at different levels. This gives the improved control synthesis problem as follows,
@@
where $\xi$ denotes the relaxation variable for soft risk threshold $r_{s}$ and $K\!>\!0$ is a constant to balance the satisfaction of the co-safety formulas and risks of safety formulas. The risk metric is minimized when exceeding the soft threshold is unavoidable. In addition, $r_{s}$ filters out overly conservative behavior and $r_{h}$ ensures critical safety.
@@ V-A Experimental Study
We validate the improved framework in three scenarios, using default parameters: risk thresholds $r_{s}\!=\!1$ and $r_{h}\!=\!2$, and discounting factor $\gamma\!=\!0.8$. The video demonstration is accessible at link  [^1] https://youtu.be/r5kEMW8oPQE and the code is provided at link  [^2] https://github.com/Miracle-qi/Risk_LTL_Planning.
1) Pedestrian crossing scenario: The improved method can generate safe behavior with tunable conservatism under the same settings as in Sec. IV-C, encompassing functionalities of the original one in Eq. (13).
2) Unexpected construction scenario: Real-world traffic often contains unexpected situations that make the original specifications infeasible [3], such as traffic participants violating rules or disruptions of traffic environments. We tested the proposed methods in such a scenario where an unexpected construction zone appears in the forward lane of the ego car, as illustrated in Fig. 7. In this case, the ego vehicle must deviate from a part of the original safety specifications to reach its target. The specification of this scenario for the ego vehicle is ($\square(\neg c\wedge\neg o\wedge\neg s)\wedge\lozenge(t)$, where $c$, $o$, $s$, and $t$ are labels for construction areas, opposite lanes, sidewalks, and target areas. We assign cost values of 555, 333, and 111 to the construction area, sidewalk, and opposite lane, respectively. Figure 7 shows the generated trajectory, where the ego vehicle autonomously chooses to bypass the construction zone via the opposite lane with minimal violation and finally reaches the target. This result demonstrates that the improved control method effectively balances different risks and minimizes violations. This advantage arises from incorporating severity into LTL specifications.
@@
3) Unprotected turn scenario: We consider an intersection scenario that includes a traffic light and dynamic opponent vehicles, as depicted in Fig. 1. The ego vehicle (blue car) is expected to follow traffic signals and avoid both non-drivable areas and the opponent vehicles (red car). The specification has been introduced in the study case of Sec. II-B. An MC model is designed to represent the switching of traffic lights and the movement of the opponent vehicle. For simplicity, we assume the opponent vehicle always moves forward. The collision risk with the opponent vehicle is evaluated by calculating the overlap between risk fields of vehicles, following the method described in [14]. With the policy synthesized by the LP problem, the ego vehicle will wait at the boundary of the intersection until the light turns green, adjust its decisions based on the real-time positions of the opponent vehicles, and finally reach the target position. The whole process is illustrated in Fig. 1 and the supplementary video. This case accounts for traffic lights, non-drivable areas, and dynamic vehicles, which are typical factors in traffic. Thus, this method can handle complex situations and has the potential to scale to various scenarios.
To examine the influence of discounting factors, the corresponding risk curves are shown in Fig. 8. While the generated behavior remains identical across different discounting factors, the risk curves in the process are different. Compared to $\gamma\!=\!0.8$, the discounting factor $\gamma\!=\!0.5$ places more focus on recent events, causing the risk curves of $\gamma\!=\!0.5$ to gradually increase as the two vehicles approach each other. In contrast, the risk curves for $\gamma\!=\!0.8$ remain flatter because the risks are perceived earlier. Since placing more focus on recent events lowers perceived risk values, the thresholds for lower discounting factors are tuned down accordingly.
@@
Table I presents the maximum and average risk values with the default discounting and threshold, along with the state number and computation time for the LP problem in (13). The results show that the improved LP problem effectively manages different types of risk and maintains gross risk value within the predefined level throughout the execution process. However, computation time increases significantly as the number of states grows, making finer abstractions for realistic traffic scenarios computationally expensive.
@@
TABLE I: Risk value and computation time of three scenarios.
@@ VI Conclusion
In this paper, we extend a human-like risk metric to account for complex traffic scenarios by integrating LTL specifications. At the same time, the proposed metric makes temporal logic more practical for traffic tasks by incorporating human-like awareness. Furthermore, a risk-aware control method is improved to balance different types of risks in a human-like manner. Extensive simulation results validate the effectiveness and potential of this framework. To better handle realistic traffic, future work should incorporate strategic interactions between vehicles, reduce computational costs for finer abstractions, and learn parameters from real-road data instead of manual configuration.